OpenStudy (anonymous):

How do I simplify the radical expression 3 square root 200 + 2 square root 8

OpenStudy (anonymous):

\[3\sqrt{200}+2\sqrt{8}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3*\sqrt{2}*\sqrt{100}+2*\sqrt{4}*\sqrt{2}\]

OpenStudy (zzr0ck3r):

\[3* \sqrt{200}+2*\sqrt{8}=3*\sqrt{100}\sqrt{2}+2\sqrt{4}*\sqrt{2}\]

OpenStudy (zzr0ck3r):

30sqrt(2) + 4sqrt(2) = 34sqrt(2)

OpenStudy (anonymous):

\[34\sqrt{2}\]

OpenStudy (zzr0ck3r):

pretty odd we did it exactly the same @Krishnadas

OpenStudy (zzr0ck3r):

great minds......

OpenStudy (anonymous):

thats what we do everytime..this is the first method that comes to mind

OpenStudy (zzr0ck3r):

yeah still odd...we use same notation exactly almost.:)

OpenStudy (anonymous):

lol, okay, thanks!